OpenStudy (anonymous):

find the zeros of each function n state the multiplicity of multiple zeros y=2x^2-2x

OpenStudy (anonymous):

2x^2-2x = 0 2x(x - 1) = 0

OpenStudy (anonymous):

can you solve this?

OpenStudy (anonymous):

how do u do that

OpenStudy (anonymous):

where you have the product of 2 function = 0 either of the functions must be = 0 so either 2x = 0 or x-1 = 0 giving you 2 values of x

OpenStudy (anonymous):

so the zeros would be 1/2 n 1

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

x - 1 = 0 so x = 1 2x = 0 x = ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

is that right